# SPA: Efficient User-Preference Alignment against Uncertainty in Medical Image Segmentation

### v1

@@ 3.3 Preference Adaption with Human Feedback
As shown in the lower part of Fig. 2, $\mathbf{\hat{y}_{app}}^{(j)}$, our model also provides $K$ correction proposals $\{\mathbf{\Delta_{r_{k}}}^{(j)}\}_{k=1}^{K}$ for further selection. These correction proposals reflect the distinctive possible ways for adapting the prediction to the desired preference. To create $\{\mathbf{\Delta_{r_{k}}}^{(j)}\}_{k=1}^{K}$, K-means clustering is applied to the predictions $\{\mathbf{\hat{y}_{n}}^{(j)}\}_{n=1}^{N}$. Each cluster produces a representative mask $\mathbf{r_{k}}^{(j)}\in\{{\mathbf{r_{k}}}^{(j)}\}_{k=1}^{K}$ through inner-cluster majority vote. The pixel-wise differences from $\mathbf{\hat{y}_{app}}^{(j)}$ highlight areas for correction, forming the $K$ correction proposals $\{\mathbf{\Delta_{r_{k}}}^{(j)}\}_{k=1}^{K}$. Once the clinician selects a preferred correction proposal $\mathbf{\Delta_{r_{p}}}^{(j)}$, the corresponding representative mask $\mathbf{r_{p}}^{(j)}$ is compared with $\mathbf{\hat{y}_{app}}^{(j)}$ (both in their pre-binarized forms) to identify the locations with the greatest divergence, denoted as $P_{diff}=\{(h,w)\},h\in[0,H-1],w\in[0,W-1]$. A random location (point prompt) sampled from $P_{diff}$, along with its semantic label, is then chosen as the feedback signal. This signal is mapped into a $L$-dimensional vector interaction embedding $e_{p}$ by the feedback encoder, as shown in Fig. 2. The feedback signal for the input image $\mathbf{x}$ is initialized by randomly selecting a pixel, as no correction proposals are available at this stage.
Using the interaction embedding $e^{(j)}_{p}$, the preference distribution $p_{\theta}^{(j)}(z)$’s parameters $\theta=\{(\mu^{(j)}_{m},{\sigma^{2}_{m}}^{(j)},\pi^{(j)}_{m})\}_{m=1}^{M}$ are updated, resulting preference distribution $p_{\theta}^{(j+1)}(z)$. This adjustment is done through adaptive blocks consisting of six fully connected layers and ReLU activations, resulting in a updated preference distribution $p_{\theta}^{(j+1)}(z)$, as shown in the lower part of Fig. 2. In this step, the overall framework $f$ adapts its output to the specific human preference through interactions, expressed as: $f^{(j)}(\mathbf{x},\mathbf{\Delta_{r_{p}}}^{(j)})=f^{(j)}(\mathbf{x},e^{(j)}_{p})$. This adaptive update process allows the preference distribution $p_{\theta}^{(j+1)}(z)$ to progressively align with human preferences through sufficient interactions, as described in Section 3.4.
3.4 Preference Distribution
Our approach theoretically ensures that SPA gradually adapts to each clinician’s clinical contexts, aligning with personalized preferences through iterative feedback. We assume that each clinician’s interactions, based on their preference and stemming from different medical contexts, are independent and identically distributed (i.i.d.). To represent all possible user preferences for different clinical needs, we model the preference distribution $p_{\theta}(z)$ as a mixture of $M$ Gaussian distributions:
where $\mathcal{N}(z\mid\mu_{m},\sigma_{m}^{2})$ denotes the Gaussian component for clinician $m$, with mean $\mu_{m}$ and variance $\sigma_{m}^{2}$. The mixture weights $\pi_{m}$ satisfy $\pi_{m}>0$ and $\sum_{m=1}^{M}\pi_{m}=1$.
Each component in this mixture model reflects a clinician’s specific interaction behavior, allowing the framework to account for multiple plausible annotations. Through iterative interactions, SPA adapts this preference distribution $p_{\theta}(z)$, aligning it with the clinician’s unique segmentation preferences and adapting to their personalized clinical needs. This process ensures that the updated distribution $p_{\theta}^{(j+1)}(z)$ effectively captures clinician preferences, thus improving future predictions $\mathbf{\hat{y}_{app}}^{(j+1)}$. We provide a theoretical proof in Appendix demonstrating that, given sufficient interactions from a specific clinician $p$, the preference distribution $p_{\theta}(z)$ could converge to the clinician’s personalized distribution $\mathcal{N}(\mu_{p},\sigma_{p}^{2})$ in theory. It theoretically guarantees that SPA can ultimately reflect each clinician’s requirements, aligning closely with diverse clinical contexts and user preferences through multi-round feedback alignment.
@@ 3.5 Training Details
Our training process simulating the actual inference, by constructing interactive segmentation tasks for each batch, where Preference-aware Segmentation and Preference Adaption with Human Feedback stages alternates.
Specifically, for each training iteration, at the Preference-aware Segmentation stage, parameters of image encoder and mask decoder, denoted as $\theta^{(j)}_{PSeg}$ are updated using the cross-entropy loss: $\theta^{(j+1)}_{PSeg}=\theta^{(j)}_{PSeg}+\alpha\nabla_{\theta_{PSeg}}\mathcal{L}_{ce}(\mathbf{\hat{y}_{app}}^{(j)},\mathbf{y})$. To capture image uncertainty, we generate $N=48$ predictions by sampling from the preference distribution $p^{(j)}_{\theta}(z)$, which is modeled with of $M=16$ Gaussian components. The means, variances, and mixture weights $\{(\mu_{m}^{(0)},{\sigma_{m}^{2}}^{(0)},\pi_{m}^{(0)})\}_{m=1}^{16}$ are uniformly initialized. Additionally, we generate $K=4$ correction proposals to allow clinicians for multi-choice selection.
For the Preference Adaption with Human Feedback step, all parameters from the Preference-aware Segmentation step remain frozen. A new prediction $\mathbf{\hat{y}_{app}}^{(j+1)}$ is generated by sampling from the updated preference distribution $p^{(j+1)}_{\theta}(z)$. We denote all parameters in feedback encoder and Gaussian parameters $\{(\mu_{m}^{(j)},{\sigma_{m}^{2}}^{(j)})\}_{m=1}^{M}$ as $\theta^{(j)}_{PAF}$. $\theta^{(j)}_{PAF}$ are updated using the cross-entropy loss: $\theta^{(j+1)}_{PAF}=\theta^{(j)}_{PAF}+\alpha\nabla_{\theta_{PAF}}\mathcal{L}_{ce}(\mathbf{\hat{y}_{app}}^{(j+1)},\mathbf{y})$. Additionally, the Gaussian mixture weights $\{(\pi_{m}^{(j)})\}_{m=1}^{M}$, denoted as $\theta^{(j)}_{\pi}$, are updated with a mean squared error loss $\mathcal{L}_{mse}$ conditioned on $\mathbf{\hat{y}_{app}}^{(j+1)}$ and $\mathbf{y}$, respectively: $\theta^{(j+1)}_{\pi}=\theta^{(j)}_{\pi}+\alpha\nabla_{\theta_{\pi}}\mathcal{L}_{mse}(P(I=i\mid\mathbf{\hat{y}_{app}}^{(j+1)},P(I=i\mid\mathbf{y}))$, where $I$ stands for the Gaussian component in the preference distribution.
Table 1: SPA Outperforms SOTA in Dice Score (%). Performance analysis comparing Dice Scores across deterministic, uncertainty-aware, and interactive models. SAM-series models use clicks for interaction, while SAM-U uses bounding boxes. SPA, with its multi-choice correction proposal setting, consistently outperforms all models across diverse datasets. 1-Iter and 3-Iter indicate performance after one and three iterations, respectively.
@@ 4 Experiment
We conducted extensive experiments to validate the effectiveness of SPA across seven uncertainty segmentation tasks represented by multi-clinician, using data from various imaging modalities, including color fundus images, CT, and MRI scans. SPA consistently achieves SOTA performance compared to deterministic, uncertainty-aware, and interactive models. Notably, SPA outperforms interactive models with significantly fewer iterations and it demonstrates strong generalization on unseen clinicians. Human evaluation further shows that SPA requires less time and fewer interactions to meet human expectations. Our results also show that predictions increasingly align with selected clinicians while diverging from those excluded. Ablation studies highlight the necessity of each component for optimal segmentation. Additional analysis of correction proposal similarity across iterations and prediction changes after each interaction is detailed in Appendix.
@@ 4.1 Dataset
REFUGE2 benchmark [10] is a publicly available fundus image dataset for glaucoma analysis, including the optic cup segmentation. It includes annotations from seven independent ophthalmologists, each with an average of eight years of experience. The dataset contains 400 training images and 400 testing images.
LIDC-IDRI benchmark [1, 8] originally consists of 3D lung CT scans with semantic segmentations of possible lung abnormalities, annotated by four radiologists. We use a pre-processed version from [21], which includes 15,096 2D CT images. After an 80-20 train-test split, the training and testing datasets contain 12,077 and 3,019 images, respectively.
QUBIQ benchmark [23] investigates inter-clinician variability in medical image segmentation tasks. It includes one MRI brain tumor task (three annotations, 28 training cases, 4 testing cases); two MRI prostate tasks (six annotations, 48 training cases, 7 testing cases); one MRI brain-growth task (seven annotations, 34 training cases, 5 testing cases); and one CT kidney task (three annotations, 20 training cases, 4 testing cases).
4.2 Experimental Setup
Our network was implemented using the PyTorch platform and trained/tested on an RTX A4000 with 16GB of memory. During training, we used the Adam optimizer with an initial learning rate of $1e^{-4}$ and adjusted it using StepLR strategy. To ensure a fair comparison, deterministic methods with multiple annotations were trained using majority vote. For the SAM-series interactive models, click or bounding box prompts were randomly generated based on the original model settings. During testing, a random set of annotations was selected, fused, and binarized. This fused binary segmentation was then used as the ground truth for evaluation. Further details are provided in Appendix. We consistently used ViT/B as the backbone when vision transformer was involved in the models.
4.3 Experimental Result
4.3.1 Performance Analysis with SOTA Methods
@@
SPA consistently outperforms all methods, achieving an average Dice Score of 89.68% after three iterations: Table 1 provides a quantitative comparison of Dice Scores across datasets. The improvement is particularly notable in the LIDC segmentation task, where SPA exceeds current SOTA methods by 20%. Even after one iteration, SPA demonstrates superior performance than SOTA methods, highlighting its potential to generate high-quality segmentation with minimal interaction. Fig. 3 shows visual comparisons between SPA and other SOTA methods, presenting segmentations after six iterations for the interactive models. The results indicate that the segmentations predicted by SPA align more closely with the ground truth, especially at boundary regions.
@@
Figure 4: SPA Demonstrates Extraordinary Efficiency. Efficiency analysis comparing the average number of iterations required to reach specific Dice Scores across interactive models. Models that failed to reach the target Dice Score within six iterations are assigned an iteration count of ten. SPA consistently requires fewer iterations to achieve high-performance segmentation results.
Table 2: SPA Demonstrates Strong Prediction Alignment with Individual Clinicians. Alignment analysis by Dice Score (%) on the REFUGE2 optic cup segmentation task. The ground truth combines annotations from Clinicians 2, 4, 5, 6, and 7, excluding Clinicians 1 and 3. Predictions are compared with individual clinicians (columns) across interaction iterations (rows). The predictions show increasing alignment with included clinicians and divergence from excluded clinicians.
@@
4.3.2 Efficiency Analysis on Different Interactive Models
We conducted an efficiency analysis to quantify the number of iterations required to reach specific Dice Scores, with a maximum of six iterations. Models that failed to reach the target Dice Score within the limit were assigned an iteration count of ten. Our proposed method, SPA, consistently required fewer iterations to achieve superior segmentation performance compared to other models.
SPA outperforms all other interactive models in terms of interaction efficiency, as shown in Fig. 4. On the REFUGE dataset, it requires fewer than five iterations to achieve a Dice Score of 87%, whereas SAM-U V2 needs 5.6 iterations to reach a Dice Score of just 84%. In the LIDC dataset, SPA achieves 75% Dice Score with only 2.7 iterations on average, while other models typically require four or more iterations. Similarly, on the Kidney dataset, SPA reaches a Dice Score of 92% with 3.2 iterations, while other SAM-series models often need four to ten iterations. Notably, SPA reaches a Dice Score of 95% within eight iterations, but other models struggle to converge within the ten-iteration limit.
SPA demonstrates generalization and robustness to unseen preferences. To show this, we left one clinician’s annotations entirely out of the training phase and evaluated on the REFUGE2 optic cup segmentation task. SPA achieves a 75% Dice Score in 5.63 iterations on average, with other models requiring more than six iterations to reach similar accuracy. For a Dice Score of 84%, SPA requires an average of 7.69 iterations, surpassing the second-best model, SAM-U V2, by 0.64 iterations.
SPA’s improved efficiency is further substantiated with human user study. This user study is based on the REFUGE dataset, conducted by two PhD students with medical expertise. One individual required an average of 6.62 seconds and 4.33 iterations per image with MedSAM, but only 4.34 seconds and 2.46 iterations with SPA. Similarly, the other individual took 7.40 seconds and 5.60 iterations per image with MedSAM, but just 4.77 seconds and 3.58 iterations with SPA. This significant reduction in both time and interactions demonstrates SPA’s superior efficiency.
@@ 4.3.3 Prediction Alignment with Clinician Feedback in Model Adaptation
Table 2 shows prediction alignment results for optic cup segmentation on the REFUGE2 dataset, comparing SPA’s predictions with individual clinicians’ annotations after each iteration. In this case, the ground truth is defined as the weighted average of annotations from Clinicians 2, 4, 5, 6, and 7, excluding Clinicians 1 and 3. After each interaction, the Dice Scores for the included clinicians consistently improve, indicating that SPA is adapting to the desired context. For example, Clinician 2’s score increases from 78.92% to 81.54% (+3.32%), with similar positive trends for Clinicians 4, 5, 6, and 7 (+ 1.25%, +2.28%, +1.38%, and +3.35%, respectively). In contrast, the Dice Scores for the excluded clinicians, such as Clinicians 1 and 3, show a consistent decline (-2.49% and -0.81%), indicating that the model is effectively excluding irrelevant clinicians. This multi-choice-based refinement demonstrates SPA’s ability to align predictions with the included clinicians while excluding those not part of the ground truth. Visualization of this alignment are provided in Appendix.

### v2

@@ 3.4 Preference Adaption with Human Feedback
Receiving user’s choice as preference. Beyond $\mathbf{y_{app}}^{(j)}$, our model also provides $K$ Representative Segmentation Candidates $\{\mathbf{r_{k}}^{(j)}\}_{k=1}^{K}$ for further selection, as shown in the lower part of Fig. 2. These candidates reflect distinct possible ways to adapt the prediction. To create $\{\mathbf{r_{k}}^{(j)}\}_{k=1}^{K}$, K-means clustering is applied to the predictions $\{\mathbf{{y}_{n}}^{(j)}\}_{n=1}^{N}$. Each cluster produces a representative segmentation candidate $\mathbf{r_{k}}^{(j)}\in\{{\mathbf{r_{k}}}^{(j)}\}_{k=1}^{K}$ as the centroid of the cluster. We then highlight the pixel-wise differences $\{\mathbf{\Delta_{r_{k}}}^{(j)}\}_{k=1}^{K}$ between $\mathbf{{y}_{app}}^{(j)}$ and $\mathbf{r_{k}}^{(j)}$ to the users. The positive and negative differences are represented in different colors for the convenience of selection. The users then pick their preferred $\mathbf{\Delta_{r_{u}}}^{(j)}$, which determines the corresponding representative segmentation candidate $\mathbf{r_{u}}^{(j)}$. The $\mathbf{r_{u}}^{(j)}$ is then sent into the interaction encoder mapping into a $L$-dimensional feedback embedding $e_{u}$ (Fig. 2). We use SAM prompt encoder, including its point and mask branches, as our interaction encoder, allowing it to accept both point and mask inputs.
Updating preference distribution. Given the interaction embedding ${e^{(j)}_{u}}$, the preference distribution $p_{\theta}^{(j)}(z)$ will be updated to $p_{\theta}^{(j+1)}(z)$ with parameters $\theta=\{(\mu^{(j+1)}_{m},{\sigma^{2}_{m}}^{(j+1)},\pi^{(j+1)}_{m})\}_{m=1}^{M}$. Specifically, since obtaining ground-truth embedding samples directly from a given mask is challenging, we replace the analytically updated GMMs with a neural network-based predictor, that we call the Preference Distribution Generator. It adaptively constructs GMMs by predicting their parameters $\theta$ based on human feedback $e^{(j)}_{u}$. This approach employs amortized inference to efficiently estimate GMM parameters. Given $e^{(j)}_{u}$, we use six forward layers to jointly predict $\mu^{(j)}_{m}$ and ${\sigma_{m}}^{(j)}$ from a $2m$-length vector, and use another six forward layers to predict $\pi^{(j)}_{m}$ (Fig. 2 lower part). The predictor of $\mu^{(j)}_{m}$ and ${\sigma_{m}}^{(j)}$ is updated by the backpropagation from the final supervision $y$, while the predictor of $\pi^{(j)}_{m}$ is separately updated by the supervision from ground-truth $\pi^{GT(j)}_{m}$, since given $\mu^{(j)}_{m}$ , ${\sigma_{m}}^{(j)}$, and $\mathbf{y}$, we can get the analytical solution for $\pi^{(j)}_{m}$ from the GMM E-step (Algorithm 1).
Segmentations with iteratively aligned preference. In the next iteration, new latent variables $\{\mathbf{z_{n}}^{(j+1)}\}_{n=1}^{N}$ are sampled from the updated preference distribution $p_{\theta}^{(j+1)}(z)$ and undergo Preference-aware Segmentation together with the input image $\mathbf{x}$ and interaction embedding $e_{u}^{(j)}$.
Inference across input images. For each input image $\mathbf{x}$, $r_{u}$ is initialized as a user point prompt (similar to SAM), and an initial $p_{\theta}(z)$ is predicted based on $r_{u}$.
@@ 3.5 Training Details
In training, for each $\mathbf{x}$, we initialize the ground truth $\mathbf{y}$ as a stochastic combination of collected multi-rater labels: $\mathbf{y}={\mathds{1}}_{0.5}\left(\sum{A\times Y}\right)$, where $A$ represents the weights uniformly sampled from $[0,1]$ and $Y$ denotes the set of user annotations. We then simulate user interaction and iteratively update the distribution generator to produce $p_{\theta}^{(j)}(z)$ from $r_{u}$, and the segmentation model to predict $\{\mathbf{y_{n}}^{(j)}\}_{n=1}^{N}$ based on $p_{\theta}^{(j)}(z)$. The initial $\mathbf{r_{u}}$ is simulated by randomly selecting a point from the ground truth. Supervision is then applied to ensure that the final jointly produced aggregation $\mathbf{y_{app}}^{(j)}$ matches ground truth $\mathbf{y}$. The full training pipeline is in Algorithm 1.
Table 1: SPA Outperforms the SOTA in Dice Score (%). We compared deterministic, uncertainty-aware, and interactive models with Dice Score as the metric. SAM-series models use clicks for interaction, while SAM-U uses bounding boxes. SPA, with its multi-choice representative segmentation candidate mechanism, consistently outperforms the other models for the seven tasks. 1-Iter and 3-Iter indicate performance after one and three iterations, respectively.
@@ 4 Experiment
We conducted experiments to validate the effectiveness of SPA across seven uncertainty segmentation tasks represented by multi-clinician, using data from three medical imaging modalities: color fundus images, lung lesion and kidney CT scans, brain and prostate MRI scans. SPA consistently achieves SOTA performance compared to deterministic, uncertainty-aware, and interactive models. Notably, SPA outperforms interactive models with significantly fewer iterations and it demonstrates strong generalization on unseen users. Human evaluation further shows that SPA requires less time and fewer interactions to meet human expectations. Our results also show that predictions increasingly align with selected users while diverging from those excluded. Ablation studies highlight the necessity of each component for optimal segmentation. Additional analysis of representative segmentation candidates’ similarities across iterations and prediction changes after each interaction are detailed in the appendix.
@@ 4.1 Dataset
REFUGE2 benchmark [10] is a publicly available fundus image dataset for glaucoma analysis, including optic cup segmentation. REFUGE2 includes annotations from seven ophthalmologists, each with an average of eight years of experience. It contains 400 training and 400 test images.
LIDC-IDRI benchmark [1, 7] originally consisted of 3D lung CT scans with semantic segmentations of possible lung abnormalities, annotated by four radiologists. We use a pre-processed version from [21] of 15,096 2D CT images. After an 80-20 train-test split, the training and test datasets contain 12,077 and 3,019 images, respectively.
QUBIQ benchmark [23] consists of one MRI brain tumor task (three annotations, 28 training cases, 4 testing cases); two MRI prostate tasks (six annotations, 48 training cases, 7 testing cases); one MRI brain-growth task (seven annotations, 34 training cases, 5 testing cases); and one CT kidney task (three annotations, 20 training cases, 4 testing cases).
4.2 Implementation Details
Our network was implemented using PyTorch v1.12 and trained/tested on an RTX A4000 with 16GB of memory. During training, we used the Adam optimizer with an initial learning rate of $1e^{-4}$ and adjusted it using StepLR strategy. $L=256$ features were extracted to generate feature embeddings and $N=48$ latent variables were sampled from the preference distribution $p^{(j)}_{\theta}(z)$, which was modeled with $M=16$ Gaussian components. We generated $K=4$ representative segmentation candidates to allow multi-choice selection and set $MAX\_USER\_{ITERATIONS}$ to 6. It takes 0.2 second per iteration to generate the prediction and the representative segmentation candidates, acceptable for human-in-the-loop settings.
We used SAM-ViT/B as the segmentation backbone. Additional implementation details are provided in the appendix. Deterministic segmentation methods with multiple annotations were trained with majority vote labels. For SAM-series interactive models, click or bounding box prompts were uniformly generated based on the original model settings.
4.3 Experimental Results
4.3.1 Performance Comparison with SOTA Methods
@@
SPA consistently outperforms all methods, achieving an average Dice Score of 89.68% after three iterations. Table 1 provides a quantitative comparison using Dice Score as the metric. The improvement is particularly notable for the LIDC segmentation task, where SPA improves on current SOTA methods by 20%. Even after one iteration, SPA demonstrates superior performance relative to the other methods. Fig. 3 shows visual comparisons between SPA and other SOTA methods, presenting segmentations after six iterations for the interactive models. The results qualitatively indicate that the segmentations predicted by SPA align more closely with the ground truth, especially in the boundary regions.
@@
Figure 4: SPA Demonstrates Extraordinary Efficiency. Efficiency analysis comparing the average number of iterations required to reach specific Dice Scores across interactive models. Models that failed to reach the target Dice Score within six iterations are assigned an iteration count of ten. Failure rate is provided in the appendix. SPA consistently achieves lower failure rates and requires fewer iterations.
4.3.2 Efficiency Analysis on Different Interactive Models
We conducted an efficiency analysis to quantify the number of iterations required to reach specific Dice Scores, with a maximum of six iterations. Models that failed to reach the target Dice Score within the limit were assigned an iteration count of ten. The failure rate is provided in the appendix. Our proposed method, SPA, consistently has lower failure rates and requires fewer iterations to achieve superior segmentation performance compared to other models.
SPA outperforms all other interactive models in terms of interaction efficiency, as shown in Fig. 4. On the REFUGE2 dataset, it requires fewer than five iterations to achieve a Dice Score of 87%, whereas SAM-U V2 needs 5.6 iterations to reach a Dice Score of just 84%. In the LIDC dataset, SPA achieves 75% Dice Score with only 2.7 iterations on average, while other models typically require four or more iterations. Similarly, on the Kidney dataset, SPA reaches a Dice Score of 92% with 3.2 iterations, while other SAM-series models often need four to ten iterations. Notably, SPA reaches a Dice Score of 95% within eight iterations, but other models struggle to converge within the ten-iteration limit.
Table 2: SPA Demonstrates Strong Prediction Alignment with Individual Users. Alignment analysis by Dice Score (%) on the REFUGE2 optic cup segmentation task. The ground truth combines annotations from Clinicians 2, 4, 5, 6, and 7, excluding Clinicians 1 and 3. Predictions are compared with individual clinicians (columns) across interaction iterations (rows). The predictions show increasing alignment with included clinicians and divergence from excluded clinicians.
@@ 4.3.2 Efficiency Analysis on Different Interactive Models
SPA shows generalization and robustness to unseen preferences. To show this, we conducted a leave-one-user-out experiment three times, each time leaving a different user out during the training phase. The reported results are averaged across these three runs, while individual user statistics are provided in the appendix. On REFUGE2 optic cup segmentation task, SPA achieves a 75% Dice Score in an average of 3.01 iterations, whereas other models require more than 3.6 iterations to reach similar accuracy. For a Dice Score of 84%, SPA requires an average of 5.02 iterations, surpassing the second-best model, MedSAM, by 0.94 iterations. In addition, SPA is transferable between images for the same user. The Dice Score improves from 91.68% to 92.43% on Kidney dataset by using user-adapted GMM parameters.
SPA’s improved efficiency is further substantiated with a human user study. We invited five medical professionals, each with over five years of graduate-level expertise, to annotate 100 fundus images from the REFUGE2 dataset. Across all annotators, SPA consistently required fewer interactions and less time compared to MedSAM. For example, one individual required an average of 6.62 seconds and 4.33 iterations per image with MedSAM, but only 4.34 seconds and 2.46 iterations with SPA. Another annotator took 7.40 seconds and 5.60 iterations per image with MedSAM, but just 4.77 seconds and 3.58 iterations with SPA. Similarly, for the remaining annotators, SPA required an average of 4.56, 5.03, and 4.82 seconds per image, compared to 6.91, 8.04, and 7.60 seconds with MedSAM. This significant reduction in both time and interactions highlights SPA’s superior efficiency. User study details are provided in the appendix.
@@ 4.3.3 Prediction Alignment with Clinician Feedback in Model Adaptation
Table 2 shows prediction alignment results for optic cup segmentation on the REFUGE2 dataset, comparing SPA’s predictions with individual clinicians’ annotations after each iteration. In this case, the ground truth is defined as the weighted average of annotations from Clinicians 2, 4, 5, 6, and 7, excluding Clinicians 1 and 3. After each interaction, the Dice Scores for the included clinicians consistently improve, indicating that SPA is adapting to the desired context. For example, Clinician 2’s score increases from 78.92% to 81.54% (+3.32%), with similar positive trends for Clinicians 4, 5, 6, and 7 (+ 1.25%, +2.28%, +1.38%, and +3.35%, respectively). In contrast, the Dice Scores for the excluded clinicians, such as Clinicians 1 and 3, show a consistent decline (-2.49% and -0.81%), indicating that the model is effectively excluding irrelevant clinicians. This multi-choice-based refinement demonstrates SPA’s ability to align predictions with the included clinicians while excluding those not part of the ground truth. Visualization of this alignment is provided in the appendix.